OpenStudy (anonymous):

Write the equation of the quadratic function with roots -10 and -8 and a vertex at (-9, -3). Enter your answer as a function beginning with f(x). Do not write any spaces. Use ^ for exponents.Leave your answer in the form:f(x)=(a)(x^2+bx+c) sample answer:f(x)=(-1/8)(x^2+5x-12)

OpenStudy (amistre64):

this looks familiar :/

OpenStudy (amistre64):

you can just fill in the vertex form and scale it maybe y = a(x-h)^2 + k and determine a such that 0 = a(x-h)^2 + k -k/(x-h)^2 = a

OpenStudy (amistre64):

then expand it out to get it to a format you want